# Spatial-Network Treatment Effects: A Continuous Functional Approach

### v3

@@ 5 Conclusion
This paper develops a continuous functional framework for treatment effects that propagate through geographic space and economic networks. The theoretical contribution derives a master equation from three independent foundations—heterogeneous agent aggregation, market equilibrium, and cost minimization—with key results including the Feynman-Kac representation and the mutual information characterization of spatial-network interaction.
The methodological contribution establishes that the framework nests the no-spillover case as a testable restriction, creating a one-sided risk profile: correct inference regardless of whether spillovers exist. Monte Carlo evidence with both binary treatment timing and continuous treatment intensity confirms that conventional methods exhibit substantial bias (25–38%) when spillovers are present, while our framework maintains correct inference across all configurations.
The empirical contribution, applying the framework to minimum wage policy, finds total effects four times larger than direct effects, demonstrating the quantitative importance of spillover channels. Structural parameter estimates have economic interpretations consistent with labor market mobility and supply chain adjustment.
@@ 5.1 Limitations
The framework assumes continuous treatment functionals, which may not apply to perfectly discrete interventions. The identification strategy requires geographic and network variation; uniformly implemented policies remain unidentified. The entropy-based diagnostics require additional validation across settings.
@@ 5.2 Extensions
Several directions appear promising:
Heterogeneous agents: Full development of type-specific parameters $(\nu_{s}(\theta),\nu_{n}(\theta),\kappa(\theta))$ would enable distributional analysis.
Optimal policy: Characterizing the optimal source $S^{*}(\mathbf{x},t,\alpha)$ would provide normative guidance.
Alternative applications: The framework applies to financial contagion, technology diffusion, disease transmission, and trade policy—settings where spatial-network propagation matters.

### v4

@@ 4 Conclusion
This paper develops a continuous functional framework for treatment effects that propagate through both geographic space and economic networks. The framework bridges three economic foundations—heterogeneous agent aggregation, market equilibrium dynamics, and cost minimization—each arriving independently at the same master equation governing treatment propagation. This convergence establishes that the framework captures fundamental economic mechanisms rather than imposing ad hoc functional forms. The Feynman-Kac representation connects the partial differential equation structure to economic intuition by characterizing treatment effects as accumulated policy exposure along stochastic paths representing economic agents migrating, firms adjusting supply chains, and prices equilibrating across connected markets.
Two key theoretical results distinguish the framework. First, the spatial-network interaction coefficient equals the mutual information between geographic and network coordinates, providing a parameter-free measure of how strongly the two propagation channels reinforce each other. Second, the framework nests the no-spillover case as a testable restriction, creating a one-sided risk profile where correct inference is maintained regardless of whether spillovers exist. Monte Carlo evidence confirms that conventional estimators exhibit 25–38% bias when spillovers of empirically relevant magnitude are present, while our estimator achieves correct inference across all configurations including the no-spillover case. Event study simulations demonstrate that when spillovers are absent, all estimators track the true dynamic treatment path; when spillovers are present, only the full framework recovers the correct treatment effects.
4.1 Broader Implications
The findings have implications for both policy evaluation and econometric practice. For policy evaluation, the Monte Carlo results suggest that standard cost-benefit analyses substantially understate total impacts when spatial and network spillovers are operative. When a policy is implemented in one jurisdiction, it affects not only units within that jurisdiction directly but also units in neighboring regions through market integration and units in economically connected sectors through supply chain linkages. Policies evaluated without accounting for these channels will appear less effective than they actually are, potentially leading to suboptimal policy choices.
The framework applies to diverse settings where spatial proximity and economic networks jointly determine treatment propagation. Financial regulations transmit through both geographic banking markets and interbank lending networks; trade policies propagate through both border regions and global supply chains; technology adoption spreads through both local demonstration effects and industry knowledge networks; minimum wage policies transmit through both labor market competition and input-output linkages; disease transmission follows both geographic proximity and social contact networks. In each setting, the interaction between spatial and network channels creates amplification effects that additive specifications miss entirely.
For econometric practice, the results highlight the importance of testing rather than assuming SUTVA. The nested structure of our framework enables researchers to test whether spillovers are empirically present using standard hypothesis tests. When tests fail to reject the no-spillover null, conventional methods are appropriate and our framework adds no value. When tests reject the null, our framework is necessary to avoid substantial bias in treatment effect estimates. The one-sided risk profile—correct inference regardless of whether spillovers exist—makes the framework a robust choice when uncertainty exists about the validity of SUTVA.
The mutual information characterization of spatial-network interaction provides practical guidance for when interaction effects are likely to be quantitatively important. When industries cluster geographically, creating strong correlation between spatial coordinates and market positions, the interaction coefficient will be large and ignoring it will cause substantial omitted variable bias. When industries are geographically dispersed, the interaction may be negligible and additive specifications sufficient. Researchers can compute sample mutual information from observed location-industry patterns to assess whether interaction effects warrant explicit modeling.
@@ 4.2 Limitations
Several limitations warrant acknowledgment. First, the continuous functional representation requires treatment intensity to vary smoothly across space, time, and network position. Perfectly discrete interventions—binary treatment applied uniformly within jurisdictions—do not generate sufficient variation to identify the continuous functional. However, most policies exhibit continuous variation either in treatment intensity (policies vary by jurisdiction and time) or in treatment propensity (probability of adoption varies continuously with covariates), making the framework applicable to a broad class of empirical settings.
Second, identification requires variation in both geographic proximity and network connections. Uniformly implemented national policies without cross-sectional or temporal variation in treatment intensity cannot be evaluated using our methods. Similarly, settings where all units are equidistant or network connections are homogeneous provide insufficient variation for identification. The framework is best suited to policies with geographic boundaries or differential adoption patterns creating exogenous variation in exposure.
Third, the entropy-based moment conditions require sufficient data to estimate mutual information nonparametrically. Small samples or sparse networks may not provide reliable entropy estimates, limiting the applicability of this identification strategy. In such cases, researchers can rely on spatial regression discontinuity and network instrumental variables alone, though at some cost in efficiency. Future work developing parametric entropy estimators tailored to economic applications would expand the practical applicability of this approach.
Fourth, the framework assumes that spillovers operate through spatial diffusion and network connections captured by the coordinates $(\mathbf{x},\alpha)$. Alternative spillover mechanisms—general equilibrium effects through aggregate price changes, anticipation effects from policy announcements, or behavioral responses to peer outcomes—may operate through channels not captured by the coordinate system. Extending the framework to accommodate these additional mechanisms represents an important direction for future research.
@@ 4.3 Extensions
Several promising extensions would enhance the framework’s scope and applicability. First, developing type-specific propagation parameters $(\nu_{s}(\theta),\nu_{n}(\theta),\kappa(\theta))$ would enable distributional analysis of how treatment effects vary across unit types. Different types of economic agents may experience larger spatial spillovers through geographic mobility while others experience larger network spillovers through industry connections. Heterogeneous adjustment speeds $\kappa(\theta)$ would characterize which types adjust rapidly versus slowly to policy shocks. Such distributional extensions would inform equity considerations in policy design.
Second, characterizing optimal policy design would provide normative guidance. Given propagation dynamics governed by the master equation, what source term $S^{*}(\mathbf{x},t,\alpha)$ maximizes social welfare? The answer depends on both the objective function and the propagation structure. For objectives emphasizing equitable outcomes, optimal policies may target spillover amplification by concentrating interventions where interaction effects are strongest. For objectives emphasizing aggregate welfare, optimal policies may exploit spatial-network structure to achieve maximum total impact with minimum direct expenditure. Solving for optimal policies requires extending the framework to incorporate welfare functions and budget constraints.
Third, dynamic extensions allowing time-varying propagation parameters $(\nu_{s}(t),\nu_{n}(t),\kappa(t))$ would capture structural changes in economic linkages. Spatial diffusion may strengthen during recessions as worker mobility increases with job displacement, or weaken during booms as labor markets tighten. Network diffusion may vary with supply chain concentration or industry composition. Identifying time variation in propagation dynamics would enhance understanding of how treatment effects evolve over business cycles or structural transformations.
Fourth, alternative applications would demonstrate the framework’s breadth. Financial contagion involves spatial propagation through regional banking markets and network propagation through interbank exposures; technology diffusion combines spatial demonstration effects with network knowledge spillovers; disease transmission follows both geographic proximity and social contact networks. Each application would require adapting the coordinate system and source specification to the particular setting while maintaining the core mathematical structure. Successful applications across diverse domains would establish the framework as a general tool for treatment effect analysis in interconnected systems.
The framework provides a foundation for future research addressing treatment effect estimation when SUTVA fails. By deriving propagation dynamics from economic primitives rather than assuming ad hoc specifications, by nesting conventional no-spillover methods as testable special cases, and by providing identification strategies combining spatial discontinuities, network instruments, and entropy-based restrictions, the framework opens new possibilities for credible treatment effect estimation in the interconnected economies that characterize modern policy environments.
Acknowledgement
This research was supported by a grant-in-aid from Zengin Foundation for Studies on Economics and Finance.